can you show me step by step cleary??
7x-2(3+4x)=19

Answer by richard1234(7193)

We have

7x - 2(3 + 4x) = 19

The expression 2(3 + 4x) expands to 6 + 8x by the distributive property, so the equation is equivalent to

7x - (6 + 8x) = 19

7x - 6 - 8x = 19 (don't forget to distribute the negative sign throughout)

-x - 6 = 19, I'll multiply both sides by -1 to get rid of the negative signs.

x + 6 = -19

x = -25